OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the second equation is substituted into the first equation. 3x + 2y = -21 x - 3y = 4

OpenStudy (lynfran):

3x+2y=-21 ...equation 1 x-3y=4 ....equation 2 x=4+3y .....equation 3 sub equation 3 into equation 1 we get 3(4+3y) +2y = -21 ........

OpenStudy (lynfran):

can solve for the solutions from there??? @xangelaax

OpenStudy (anonymous):

these were the options bc i solved for y 3(3y + 4) + 2y = -21 3(-3y + 4) + 2y = -21 3x + 2(3y + 4) = -21 3x + 2(-3y + 4) = -21

OpenStudy (lynfran):

well if you look back at my previous post u will see the answer in ur options...

OpenStudy (anonymous):

sorry thanks so much

OpenStudy (lynfran):

now which one is it?

OpenStudy (anonymous):

first

OpenStudy (lynfran):

correct

OpenStudy (lynfran):

ur welcome

OpenStudy (anonymous):

yay thanks